OpenStudy (prettygirl_shynice):

may someone help with math i need help explaing like how i got answer i know the answer the answer but how i explaini know the answer its c but how i explain 6 to 8 and 18 to x a.20 b.22 c.24 will ward and +fan

OpenStudy (abdullahm):

Something like: \(\Large \sf \frac{6}{8}=\frac{18}{x}\) We need to solve for 'x' Do you know how to cross multiply?

OpenStudy (prettygirl_shynice):

yea

OpenStudy (usukidoll):

so when cross multiplication is being used we have 6x=18(8) once we multiply 18 x 8 we divide both sides by 6 to figure out what x is

OpenStudy (prettygirl_shynice):

okay 18 time 8 is 144

OpenStudy (abdullahm):

Correct

OpenStudy (usukidoll):

yes so now we have 6x=144 so divide 6 on both sides

OpenStudy (prettygirl_shynice):

okay hold on

OpenStudy (prettygirl_shynice):

75/3

OpenStudy (abdullahm):

That's wrong :c

OpenStudy (prettygirl_shynice):

25

OpenStudy (usukidoll):

wait wait wait... that is wrong.. it's 144/6

OpenStudy (abdullahm):

144/6 Now, 144 can be divided by 2, and 6 can be divided by 2.

OpenStudy (abdullahm):

So, what is 144/2 = ? and then we can divide by 3 to get the final answer :)

OpenStudy (usukidoll):

blah.. so when we divided on both sides we got x = 144/6 now we need to figure out what that fraction is

OpenStudy (prettygirl_shynice):

72

OpenStudy (abdullahm):

and now 72/3 = ?

OpenStudy (prettygirl_shynice):

24

OpenStudy (usukidoll):

yeah

OpenStudy (prettygirl_shynice):

thank u im happy

OpenStudy (abdullahm):

Good job!

OpenStudy (prettygirl_shynice):

thank u may u help with more but explain

OpenStudy (usukidoll):

:) ratios can be written in different ways 6 to 8 can also be written as 6:8 or \[\frac{6}{8}\] similarly with 18 to x that can be written as 18:x or \[\frac{18}{x}\]

OpenStudy (usukidoll):

so when the problem asked for 6 to 8 and 18 to x that's comparing the two ratios together and try to find out what our x is.. errrr. ummm we could've done this take the original ratio 6 to 8 and then multiply by 3 because our second ration is 18 to x Now I'm noticing that 6 to 8 is the original and 18 to x is a bigger version of that original ratio if we had multiplied 6 x 3 to gain 18 we could've multiplied 8 x 3 to get the answer a bit quicker

OpenStudy (prettygirl_shynice):

so how i explain

OpenStudy (usukidoll):

6 to 8 is a ratio... that can be re-written as either 6:8 or as a fraction \[\frac{6}{8}\] 18 to x is also a ratio that can be re-written as either 18: x or as a fraction \[\frac{18}{x}\] so we need to compare the 2 ratios and find out what the missing x value is \[\Large \sf \frac{6}{8}=\frac{18}{x}\]

OpenStudy (prettygirl_shynice):

k

OpenStudy (usukidoll):

so we can use cross multiplication  to get to this step \[6x=(18)(8)\]